A ball moving at 10 m/s makes an off-center elastic collision with another ball of equal mass that is initially at rest. The incoming ball is deflected at an angle of 30° from its original direction of motion. Find the velocity of each ball after the collision.


1
Expert's answer
2021-05-31T15:38:00-0400

"mv_0^2/2=mv^2/2+mu^2/2\to v_0^2=v^2+u^2\to\beta=90°-30°=60°"


"mv_0\sin 30°=mu\to u=v_0\sin30°=10\cdot\sin30°=5\ (m/s)"


"mv_0\cos 30°=mv\to v=v_0\cos30°=10\cdot\cos30°=8.66\ (m/s)"